Presented below is information related to Vaughn Company.

Cost

Retail

Beginning inventory $ 61,600 $107,300
Purchases (net) 120,170 180,700
Net markups 10,325
Net markdowns 26,679
Sales revenue 187,090
Compute the ending inventory at retail.
Ending inventory $

LINK TO TEXT

Compute a cost-to-retail percentage under the following conditions. (Round ratios to 2 decimal places, e.g. 78.74%)

Cost-to-retail percentage

(1) Excluding both markups and markdowns.

%
(2) Excluding markups but including markdowns.

%
(3) Excluding markdowns but including markups.

%
(4) Including both markdowns and markups.

%

LINK TO TEXT

Which of the methods in (b) above does the following?
(1) Provides the most conservative estimate of ending inventory.

Excluding Both Markups and Markdowns.Excluding Markdowns but Including MarkupsExcluding Markups but Including MarkdownsIncluding Both Markdowns and Markups

(2) Provides an approximation of lower-of-cost-or-market.

Excluding Both Markups and Markdowns.Excluding Markdowns but Including MarkupsExcluding Markups but Including MarkdownsIncluding Both Markdowns and Markups

(3) Is used in the conventional retail method.

Excluding Both Markups and Markdowns.Excluding Markdowns but Including MarkupsExcluding Markups but Including MarkdownsIncluding Both Markdowns and Markups

LINK TO TEXT

Compute ending inventory at lower-of-cost-or-market. (Round ratio to 2 decimal places, e.g. 78.74% and final answer to 0 decimal places, e.g. 6,225.)
Ending inventory $

LINK TO TEXT

Compute cost of goods sold based on (d). (Round answer to 0 decimal places, e.g. 6,225.)
Cost of goods sold $

LINK TO TEXT

Compute gross margin based on (d). (Round answer to 0 decimal places, e.g. 6,225.)
Gross margin $
Answer #1

ANSWER

Compute a cost-to-retail percentage under the following conditions

Answer 1.

Excluding both mark-ups and markdowns:

Cost = $61,600 + 120,170 = $181,770

Retail = $107,300 + 180,700 = $288,000

Cost-to-retail percentage = $181,770/288,000

Cost-to-retail percentage = 0.6311 = 63.11%

Answer 2.

Excluding mark-ups but including markdowns:

Cost = $61,600 + 120,170 = $181,770

Retail = $107,300 + 180,700 – 26,679 = $261,321

Cost-to-retail percentage = $181,770/261,321

Cost-to-retail percentage = 0.6956 = 69.56%

Answer 3.

Excluding markdowns but including markups:

Cost = $61,600 + 120,170 = $181,770

Retail = $107,300 + $180,700 + $10,325 = $298,325

Cost-to-retail percentage = $181,770/298,325

Cost-to-retail percentage = 0.6093 = 60.93%

Answer 4.

Including both markdowns and markups:

Cost = $61,600 + 120,170 = $181,770

Retail = $107,300 + $180,700 + $10,325 - $26,679 = $271,646

Cost-to-retail percentage = $181,770/271,646

Cost-to-retail percentage = 0.6691 = 66.91%

Which of the methods in (b) above does the following?

Provides the most conservative estimate of ending inventory.

Excluding Markdowns but Including Markups

Provides an approximation of lower-of-cost-or-market.

Excluding Markdowns but Including Markups

Is used in the conventional retail method.

Excluding Markdowns but Including Markups

a) Compute ending inventory at lower-of-cost-or-market.
Cost Retail
Beginning inventory $ 61,600 $              107,300
Purchases (net) $     120,170 $              180,700
Net markups $                10,325
$     120,170 $              298,325
Net markdowns $                26,679
$     120,170 $              271,646
Sales revenue $    187,090
Ending inventory at retail $                      84,556
Ending inventory=84556*120170/298325=34,060.49
b) Compute cost of goods sold based on (d)
Cost of goods sold=120170-34060.49= $ 86109.51
c) Compute gross margin based on (d)
Gross profit=187090-86109.51=$ 100,980.49

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************